[Hallc_running] [New Logentry] Hall C Status update (Dec. 20)

gaskelld at jlab.org gaskelld at jlab.org
Wed Dec 20 09:40:01 EST 2017

Logentry Text:
Yesterday we completed the SHMS delta scan with the new nominal quad values.  There was also some work done setting up the HMS and SHMS PID trigger thresholds and overnight a run was taken in the SHMS with the quads off and the 2nd collimator (in front of the bender installed).

Overnight, we ran into some issues with the SHMS and HMS rotation (Eric Sun is in the hall now troubleshooting). There will also issues with the target motion when trying to move from the 10 cm LH2 cell to the 4 cm cell. Dave Meekins came in to resolve the target motion issue. 

Today, we will spend some time taking data to gain match the quartz bars and perform a hodoscope threshold scan. Later, we will begin setting up the coincidence trigger and take H(e,e'p) data.

Beam for physics stops tomorrow morning at 7 am. However, Hall C will remain locked up for an extra hour so that some time can be taken to check out the big BPMs downstream of the target. The hall should be in Restricted Access by late morning.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3507525
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20171220/34cf2e40/attachment.html>

More information about the Hallc_running mailing list